Hannah Ortega is considering expanding her business. She plans to hire a salesperson to cover trade shows. Because of compensation, travel expenses, and booth rental, fixed costs for a trade show are expected to be $7,130. The booth will be open 23 hours during the trade show. Ms. Ortega also plans to add a new product line, ProOffice, which will cost $187 per package. She will continue to sell the existing product, EZRecords, which costs $110 per package. Ms. Garcia believes that the salesperson will spend approximately 13 hours selling EZRecords and 10 hours marketing ProOffice.

Required

  1. Determine the estimated total cost and cost per unit of each product, assuming that the salesperson is able to sell 89 units of EZRecords and 44 units of ProOffice.

  2. Determine the estimated total cost and cost per unit of each product, assuming that the salesperson is able to sell 200 units of EZRecords and 102 units of ProOffice.

EZRecords ProOffice
a. Total cost of sales
Cost per unit
b. Total cost of sales
Cost per unit
Answer #1

Fixed cost is allocated to the EZRecords and ProOffice in the ratio based on the selling hours i.e. 13:10

Allocation of fixed cost :-

EZRecords = 7,130 x 13/23 = 4,030

ProOffice = 7,130 x 10/23 = 3,100

(a)

EZRecords ProOffice
Total cost of sales (89 x 110) + 4,030 = $13,820 (44 x 187) + 3,100 = $11,328
Cost per unit 13,820/89 = $155.281 11,328/44 = $257.455

(b)

EZRecords ProOffice
Total cost of sales (200 x 110) + 4,030 = $26,030 (102 x 187) + 3,100 = $22,174
Cost per unit 26,030/200 = $130.150 22,174/102 = $217.392